OpenStudy (anonymous):

Choose the product -6p^3(3p^2 + 5p - 1) A.18p^5- 30p^4+ 6p^3 B.-18p^6- 30p^3- 6p C.18p^3+ 6p^2- 30p^4 D.-18p^6- 24p^3

OpenStudy (tkhunny):

(-6p^3)(3p^2) = -18p^5 So, A, B, C, and D and C are our of the running. Where does that leave us?

OpenStudy (anonymous):

what did you do with the +5p-1?

OpenStudy (tkhunny):

I ignored them. The first term is not among the options, so I didn't care if the other two terms were there.

OpenStudy (anonymous):

but don't you multiply the two exponets?

OpenStudy (tkhunny):

\((a^2)(a^3) = a^5 \ne a^{6}\)

OpenStudy (anonymous):

then it has to be A?

OpenStudy (tkhunny):

No, A is incorrect. The sign of the first term is incorrect. Did you copy it correctly?

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

Then there is no correct answer. "A" is compelling, but not quite there. "A" it the only one with the right number of terms and the right exponents. Still wrong.